OpenStudy (logo):

1.Why of the following is the probability of drawing a black card or a face card from a standard of deck cards? Quick Help! a. 1/2 b.19/26 c.3/26 d.8/13 2. What is the probability of rolling a 6 or a number less than 3 on a six sided number cube? a. 1/2 b.1/3 c.2/3 d.5/6

OpenStudy (logo):

@jim_thompson5910 @Nnesha @freckles @HannahC234

OpenStudy (mathstudent55):

How many black cards are there?

OpenStudy (logo):

8?

OpenStudy (mathstudent55):

No. Much more. Do you know what a standard deck of cards consists of?

OpenStudy (logo):

52

OpenStudy (mathstudent55):

Yes, a standard deck of cards has 52 cards. The 13 different face values are Ace, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, Jack, Queen, King The Jack, Queen, and King are called the face cards. Each of those 13 cards exists in 4 suits: diamonds (red), hearts (red), clubs (black), spades (black)

OpenStudy (mathstudent55):

There are 13 cards of each of 4 suits, Since 13 * 4 = 52, there are 52 cards in a deck. The 13 hearts and 13 diamonds cards are red, so there are 2 * 13 = 26 red cards.

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (mathstudent55):

There are 2 suits that are black: clubs and spades. Since each suit has 13 cards, there are 2 * 13 = 26 black cards

OpenStudy (mathstudent55):

You need to find out how many cards you are interested in drawing and compare to the total number of card sin a deck. Clubs (black): 13 cards Spades (black): 13 cards Hearts (red) face cards: 3 cards Diamonds (red) face cards: 3 cards You are interested in drawing a card from the (13 + 13 + 3 + 3) black or face cards out of a total of 52 cards in the deck.

OpenStudy (mathstudent55):

Do you understand how to answer problem 1?

OpenStudy (logo):

Oh, so 32/52. The answer is d? 8/13?

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

For problem 2, you know that a die has 6 sides, each one with a number from 1 to 6. There is an equal probability of getting any of those 6 numbers. Now count how many numbers you are interested in and divide by 6.